OpenStudy (linn99123):

NEED HELP MADEL AND FAN m/18=15/12 What is the solution to the proportion???

OpenStudy (anonymous):

Solution for m/18=15/12 equation: m in (-oo:+oo) m/18 = 15/12 // - 15/12 m/18-(15/12) = 0 m/18-5/4 = 0 1/18*m-5/4 = 0 // + 5/4 1/18*m = 5/4 // : 1/18 m = 5/4/1/18 m = 45/2 m = 45/2

OpenStudy (linn99123):

so 14.4?

OpenStudy (chelseatheruler):

22.5 omg

OpenStudy (anonymous):

Chelsea is right it's 22.5 you cut 45 in half. :)